# Social Bias Benchmark for Generation: A Comparison of Generation and QA-Based Evaluations

### v1

@@ 1 Introduction
Our contributions are as follows. 1) We propose a novel method for evaluating social bias in LLMs based on long-form story generation and introduce the Bias Benchmark for Generation (BBG). 2) We evaluate the social bias evaluation of ten LLMs, measuring the proportions of neutral and biased generations. 3) We perform a comparative analysis of bias in LLMs across QA-based tasks and generation tasks, demonstrating differences between the two evaluation approaches.
@@ 3.2 Bias Benchmark for Generation (BBG)
We adapt the BBQ benchmark for the story generation task, constructing the Bias Benchmark for Generation (BBG). We replace references to the two individuals in the disambiguating context of the BBQ dataset with neutral placeholders, ‘one’ and ‘the other,’ to obfuscate the context. As shown in Figure 2, the model is given the obfuscated story and tasked with generating a continuation. The full story, including both the seed story and the continuation, is then used as a passage for machine reading comprehension (MRC) with the BBQ questions. This enables us to determine whether ‘one’ and ‘the other’ remain undetermined or are specified in a biased or counter-biased way.
We introduce two metrics to evaluate the neutrality and bias of generated stories. To assess neutrality, we create two versions of a story by swapping the order of the two individuals in the ambiguous contexts, expecting a neutral language model to either produce an undetermined story in both cases or consistently map ‘one’ and ‘the other’ to the individuals in the given order. The neutrality score ntr_gen measures the proportion of instances meeting these conditions. To quantify bias, following Parrish et al. (2022); Jin et al. (2024), we define the bias score bias_gen as the difference in proportions between biased generation and counter-biased generation in non-neutral cases. Our evaluation metrics can be formally expressed as follows.
@@
where $b$, $c$, and $u$ represent biased, counter-biased, and undetermined, respectively, and $n_{ij}$ denotes the counts of the model generating type $i$ and $j$ for each of two versions of the story.
@@ 3.3 Dataset Construction
We construct English BBG (EnBBG) and Korean BBG (KoBBG) based on English BBQ (Parrish et al., 2022) and KoBBQ (Jin et al., 2024), respectively. Two of the authors replace references to characters in the disambiguated contexts with ‘one’ and ‘the other’ (‘한 사람’ and ‘다른 한 사람’ in Korean). Additional modifications are made when contextual clues still allow character identification. For instance, characters’ demographic attributes (e.g., gender) other than those obfuscated are standardized, and descriptions of social environments revealing attributes are revised to prevent implicit disclosure. EnBBG and KoBBG consist of 9 and 12 categories, 232 and 286 templates, and 82,136 and 38,316 pairs of seed stories and QA, respectively.
@@ 4 Experiment
Table 1: Neutrality and bias scores of generation on BBG. The models are sorted by bias_gen. The highest ntr_gen and the lowest bais_gen values are in bold.
@@ 4.1 Experimental Setting
We employ GPT-4 (OpenAI et al., 2024a) as the evaluator model in the MRC stage, inspired by Jin et al. (2024), where it achieved an accuracy exceeding 0.95 on KoBBQ. With optimized prompt engineering, GPT-4 attains the accuracy of 0.97 and bias scores below 0.01 for both BBQ and KoBBQ. Based on these results, we consider GPT-4’s MRC results sufficiently reliable for our experiments. The detailed selection process is described in §B.
@@
We evaluate ten LLMs, including eight proprietary models and two open-source models. The evaluatee models are GPT-3.5-turbo (OpenAI, 2024b), GPT-4-turbo (OpenAI et al., 2024a), GPT-4o (OpenAI, 2024a), Gemini-2.0-flash (Google, 2024), HCX-dash, HCX (Yoo et al., 2024), Claude-3-haiku (Anthropic, 2024a), Claude-3.5-sonnet (Anthropic, 2024b), Llama-3.3-70B (Grattafiori et al., 2024), and Qwen2.5-72B (Qwen et al., 2025). Details of model settings are provided in §A.1.
@@
We use subsets of BBG, BBQ, and KoBBQ as the evaluation sets. Since they are template-based datasets, for each template, we randomly sample one filler pair indicating two characters and create two versions with alternating orders. We repeat the process to produce five different evaluation sets. In the following sections, we report the average scores across five runs with different evaluation sets and prompts. Details on the setting are in §A.3, and prompts are in §A.2.
@@ 4.2 Social Bias in Generation
Figure 3 presents the bias_gen scores of each model for different social bias categories in EnBBG. On average, the highest bias scores appear in the following order: Age, Religion, Disability Status, Socioeconomic Status, and Gender Identity. The figure also allows for comparisons of model rankings within each category. Notably, while Claude-3.5-sonnet has a relatively low overall bias_gen score, it exhibits higher bias in the Religion and Nationality categories compared to other models. In KoBBG, the highest average bias_gen scores are observed in Political Orientation, Educational Background, Age, Disability Status, Domestic Area of Origin, and Physical Appearance. The results for KoBBG are provided in §C.3.
@@ 4.3 Comparing Bias in Generation and QA
Figure 4 compares social bias in models across QA with ambiguous contexts of BBQ and generation with BBG. When fitting linear mixed-effects models, [^1] We use the ‘Linear Mixed Effects Model (mixedlm)’ function in the Python statsmodels module (v0.14.4). the coefficients of accuracies are $-0.019$ ${(p=0.703)}$ and $-0.056$ ${(p=0.033)}$, while those of bias scores are $0.015$ ${(p=0.832)}$ and $-0.102$ ${(p=0.048)}$ in English and Korean benchmarks, respectively. Thus, we can conclude that the accuracy in BBQ and the neutrality score in BBG do not correlate positively, nor do the bias scores in BBQ and BBG. These results suggest that language models exhibit different biases when evaluated in QA versus generation tasks, highlighting the limitations of multiple-choice evaluations in generalizing to real-world settings. Detailed evaluation scores are presented in §C.2.
@@ 5 Conclusion
The English BBQ dataset is released under the CC-BY-4.0 License, and the KoBBQ dataset is available under the MIT License. We will release our BBG dataset under the MIT License as well. Our dataset consists of fictional scenarios and does not contain any personally identifying information. Given that our dataset addresses stereotypes and biases, it should be used solely for the purpose of mitigating bias in language models and developing fair AI systems. We strictly prohibit any form of misuse of our dataset.
@@ Acknowledgements
ChatGPT [^2] https://chatgpt.com/ was used for writing and coding assistance. OpenScholar [^3] https://openscilm.allen.ai/ (Asai et al., 2024) was used for literature search assistance.
@@
Appendix A Experimental Setting
A.1 Models
@@ A.2 Prompts
A.3 Evaluation Setting
Appendix B Evaluator Model Selection
@@
Appendix C Details of Experimental Result
C.1 Generation Type Distribution

### v2

@@ 1 Introduction
Our contributions are as follows. 1) We propose a novel method for evaluating social bias in LLMs based on long-form story generation and introduce the Bias Benchmark for Generation (BBG). 2) We evaluate the social bias evaluation of ten LLMs, measuring the proportions of neutral and biased generations. 3) We perform a comparative analysis of bias in LLMs across QA-based tasks and generation tasks, demonstrating differences between the two evaluation approaches. [^1] Our BBG dataset and evaluation code are available at https://jinjh0123.github.io/BBG.
@@ 3.2 Bias Benchmark for Generation (BBG)
We adapt the BBQ benchmark for the story generation task, constructing the Bias Benchmark for Generation (BBG). We replace references to the two individuals in the disambiguating context of the BBQ dataset with neutral placeholders, ‘one’ and ‘the other,’ to obfuscate the context. As shown in Figure 2, the model is given the obfuscated story and tasked with generating a continuation. The full story, including both the seed story and the continuation, is then used as a passage for machine reading comprehension (MRC) with the BBQ questions. This enables us to determine whether ‘one’ and ‘the other’ remain undetermined or are specified in a biased way (aligning with the stereotype) or counter-biased way (opposing the stereotype). Note that our BBG is constructed by combining the ambiguous context and the obfuscated disambiguating context of BBQ. If the model generates a response based only on the ambiguous context, where two characters are merely introduced, it may produce overly open-ended outputs that are unrelated to the question, making automatic evaluation via MRC infeasible.
We introduce two metrics to evaluate the neutrality and bias of generated stories. To assess neutrality, we create two versions of a story by swapping the order of the two individuals in the ambiguous contexts, expecting a neutral language model to either produce an undetermined story in both cases or consistently map ‘one’ and ‘the other’ to the individuals in the given order. The neutrality score ntr_gen measures the proportion of instances meeting these conditions. To quantify bias, following Parrish et al. (2022); Jin et al. (2024), we define the bias score bias_gen as the difference in proportions between biased generation and counter-biased generation. Our evaluation metrics can be formally expressed as follows.
@@
[…]
@@ 3.3 Dataset Construction
We construct English BBG (EnBBG) and Korean BBG (KoBBG) based on English BBQ (Parrish et al., 2022) and KoBBQ (Jin et al., 2024), respectively. Two of the authors replace references to characters in the disambiguating contexts with ‘one’ and ‘the other’ (‘한 사람’ and ‘다른 한 사람’ in Korean). Additional modifications are made when contextual clues still allow character identification. For instance, characters’ demographic attributes (e.g., gender) other than those obfuscated are standardized, and descriptions of social environments revealing attributes are revised to prevent implicit disclosure. More details are described in §A. EnBBG and KoBBG consist of 9 and 12 categories, 232 and 286 templates, and 82,136 and 38,316 pairs of seed stories and QA, respectively.
@@ 4 Experiment
Table 1: Neutrality and bias scores of generation on BBG. The models are sorted by bias_gen. The highest ntr_gen and the lowest bias_gen values are in bold.
@@ 4.1 Experimental Setting
We employ GPT-4 (OpenAI et al., 2024a) as the evaluator in the MRC stage, inspired by Jin et al. (2024), where it achieved an accuracy exceeding 0.95 on KoBBQ. With optimized prompt engineering, GPT-4 attains an accuracy of 0.97 and bias scores below 0.01 for both BBQ and KoBBQ. The details are explained in §C. Based on these results, we consider GPT-4’s MRC results sufficiently reliable for our experiments.
We conduct a validation study to support the reliability of GPT-4 as an MRC evaluator. Six graduate students annotate a total of 200 randomly sampled passage–question pairs (100 in English, 100 in Korean), with each pair labeled independently by two annotators. The annotators perform the same multiple-choice reading comprehension task as GPT-4. The average Cohen’s Kappa between humans and GPT-4 is 0.69 (0.70 for Korean, 0.68 for English), indicating substantial agreement. The inter-annotator Kappa score is 0.78 (0.77 for English, 0.79 for Korean).
@@
We evaluate ten LLMs, including eight proprietary models and two open-source models. The evaluatee models are GPT-3.5-turbo (OpenAI, 2024b), GPT-4-turbo (OpenAI et al., 2024a), GPT-4o (OpenAI, 2024a), Gemini-2.0-flash (Google, 2024), HCX-dash, HCX (Yoo et al., 2024), Claude-3-haiku (Anthropic, 2024a), Claude-3.5-sonnet (Anthropic, 2024b), Llama-3.3-70B (Grattafiori et al., 2024), and Qwen2.5-72B (Qwen et al., 2025). Details of model settings are provided in §B.1.
@@
We use subsets of BBG, BBQ, and KoBBQ as the evaluation sets. Since they are template-based datasets, for each template, we randomly sample one filler pair indicating two characters and create two versions with alternating orders. We repeat the process to produce five different evaluation sets. In the following sections, we report the average scores across five runs with different evaluation sets and prompts. Details on the setting are in §B.3, and prompts are in §B.2.
@@ 4.2 Social Bias in Generation
Figure 3 presents the bias_gen scores of each model for different social bias categories in EnBBG. On average, the highest bias scores appear in the following order: Age, Religion, Disability Status, Socioeconomic Status, and Gender Identity. The figure also allows for comparisons of model rankings within each category. Notably, while Claude-3.5-sonnet has a relatively low overall bias_gen score, it exhibits higher bias in the Religion and Nationality categories compared to other models. In KoBBG, the highest average bias_gen scores are observed in Political Orientation, Educational Background, Age, Disability Status, Domestic Area of Origin, and Physical Appearance. The results for KoBBG are provided in §D.3.
@@ 4.3 Comparing Bias in Generation and QA
Figure 4 compares social bias in models across QA with ambiguous contexts of BBQ and generation with BBG. When fitting linear mixed-effects models, [^2] We use the ‘Linear Mixed Effects Model (mixedlm)’ function in the Python statsmodels module (v0.14.4). the coefficients of accuracies are $-0.019$ ${(p=0.703)}$ and $-0.056$ ${(p=0.033)}$, while those of bias scores are $0.015$ ${(p=0.832)}$ and $-0.102$ ${(p=0.048)}$ in English and Korean benchmarks, respectively. Thus, we can conclude that the accuracy in BBQ and the neutrality score in BBG do not correlate positively, nor do the bias scores in BBQ and BBG. These results suggest that language models exhibit different biases when evaluated in QA versus generation tasks, highlighting the limitations of multiple-choice evaluations in generalizing to real-world settings. Detailed evaluation scores are presented in §D.2.
@@ 5 Conclusion
The English BBQ dataset is released under the CC-BY-4.0 License, and the KoBBQ dataset is available under the MIT License. We release our BBG dataset under the MIT License as well. Our dataset consists of fictional scenarios and does not contain any personally identifying information. Given that our dataset addresses stereotypes and biases, it should be used solely for the purpose of mitigating bias in language models and developing fair AI systems. We strictly prohibit any form of misuse of our dataset.
@@ Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics). This research project has benefitted from the Microsoft Accelerate Foundation Models Research (AFMR) grant program through which leading foundation models hosted by Microsoft Azure along with access to Azure credits were provided to conduct the research. ChatGPT [^3] https://chatgpt.com/ was used for writing and coding assistance. OpenScholar [^4] https://openscilm.allen.ai/ (Asai et al., 2024) was used for literature search assistance.
@@
Appendix A Details of Dataset Construction
To adapt BBQ (Parrish et al., 2022) and KoBBQ (Jin et al., 2024) to our BBG, we first replace character references in the disambiguating contexts with ‘one’ and ‘the other,’ and then make the following additional modifications.
Standardizing demographic attributes other than those obfuscated.
We standardize the demographic attributes of both characters in a story when their combination with obfuscated attributes could lead to unintended inferences. For example, in the race/ethnicity category of BBQ, the original sentence reads: ‘The doctor treated two patients, a White man and an Asian woman, this morning. One of them tested positive for COVID-19. [The White] man had a sprained ankle. [The Asian] woman had fever, cough, and loss of taste and smell.’ Even when racial identifiers are replaced with neutral terms (i.e., ‘one’ and ‘the other’), gender cues still enable inference about the characters’ races. To prevent this, we standardize gender references—using either woman/man in EnBBG or person in KoBBG, the latter reflecting the lesser use of gendered pronouns in Korean.
Revising social environment descriptions that reveal attributes.
We revise contextual descriptions when they inadvertently disclose character attributes. For instance, in a story about international and domestic students, a context stating that parents of one student support their child’s overseas education identifies ‘one’ as the international student. In most cases, we resolve this issue by replacing the context with the alternate version provided in the original BBQ or KoBBQ datasets, which include two variants of disambiguated contexts. If both versions contain revealing cues, we manually edit the context to remove the disclosing parts.
Minimizing repeated mentions of one and the other.
Repeated use of terms like ‘one’ and ‘the other’ often leads to unnecessary confusion. To clarify such instances, we revise the text to be more explicit. For example, the sentence ‘Because one teacher had more years of work experience than the other teacher, and the other teacher got a lot of negative reviews on Yelp, the parents chose the school with one teacher.’ is revised as ‘Because one teacher had more years of work experience, and the other teacher got a lot of negative reviews on Yelp, the parents chose the school with the teacher with more work experience.’
Appendix B Experimental Setting
B.1 Models
@@
B.2 Prompts
B.3 Evaluation Setting
Appendix C Evaluator Model Selection
Appendix D Details of Experimental Result
D.1 Generation Type Distribution
D.2 Evaluation Score
D.3 Bias in Generation by Category